the tens digit of a two digit number is 2 more than twice the units digits the sum of the digits is 11, find the number!
Answer by mananth(16949)

let the units digit be x
the tens digit will be 2x+2
...
2x+2 +x=11
3x+3=11
-3
3x=11-3
3x=9
/3
x= 3
tens digit = 2x+2
2*3+2
=8
the number is 83.
...
m.ananth@hotmail.ca